# On Achievable Spectral Efficiency Using Adaptive Transmission Over Terrestrial Coherent FSO Links

### v9

@@ Abstract
Terrestrial free-space optical (FSO) communication systems, while designed to operate on large unlicensed optical bandwidths, are power-constrained due to strict eye safety regulations. The channel fluctuation inherent in terrestrial FSO links also limits the received optical power. Consequently, the available signal-to-noise ratio (SNR) per Hz could become limited; this holds for future terrestrial systems based on coherent optical communications. An efficient and adaptive transmission mechanism is thus crucial at the optical transmitter. However, a critical assessment of the impact of adaptive transmission in terrestrial FSO systems has received less attention in the literature. This work studies terrestrial coherent FSO communication systems employing adaptive beam transmission while detection receiver operate under shot noise-limited conditions. Specifically, we propose a novel exact closed-form expression for the average spectral efficiency of a coherent FSO system with optimal adaptive transmissions over the gamma-gamma turbulence channel with pointing errors. More importantly, we provide a detailed assessment of the impact of turbulence and pointing error impairments on the coherent FSO system performance, revealing several novel and counterintuitive insights. In particular, the extensive numerical results help elucidate the intricacies of analyzing these terrestrial FSO systems and clarify a few misconceptions alluded to in recent related literature.
@@ I Introduction
Laser-based free-space optical (FSO) communication has gained significant attention as a viable alternative to conventional radio-frequency (RF) and fiber-optic technologies for delivering efficient backhaul connectivity in outdoor wireless networks. This growing interest is driven by several inherent advantages of FSO systems, including the availability of vast unlicensed bandwidth, ease of deployment, inherent physical-layer security, and low-cost transceiver implementation [1]. Consequently, the scope of FSO-based wireless technologies has broadened considerably, finding applications in a wide range of scenarios such as ad hoc wireless networks for tactical and emergency response situations [1], terrestrial last-mile broadband access [2], deep-space communications [3], inter-satellite links [4], and various civil and military operations [5]. In particular, the demand for ultra-high data rates in next-generation wireless networks, including 5G and 6G systems, has spurred renewed interest in coherent FSO technologies, with recent terrestrial demonstrations achieving net data rates exceeding 400 Gbps [6], [7], and 800 Gbps per channel [8].
Detection schemes in most practical FSO communication systems fall into two main categories:
@@
(i) Intensity modulation with direct detection (IM-DD): In this modulation-detection scheme, information is encoded in the intensity of the transmitted optical beam, effectively utilizing only one degree of freedom. At the receiver, the optical signal is detected directly via its impingement on a photodetector. IM-DD is favored in commercial systems due to its simplicity and low-cost implementation but suffers from limited receiver sensitivity, restricting its use to short-range terrestrial links [9], [10].
@@
(ii) Coherent heterodyne detection (HD): Coherent FSO communication systems exploit both the amplitude and phase of the optical carrier to convey information, enabling two degrees of freedom. In HD, a locally generated optical signal—typically offset in frequency from the received signal—is mixed with the incoming beam prior to photodetection. This detection approach significantly improves receiver sensitivity. Additionally, coherent receivers enable high selectivity via narrowband RF filtering, in contrast to the broader optical filtering in IM-DD based FSO communication systems [10].
Despite their numerous advantages, terrestrial FSO communication systems face two primary practical challenges. First, FSO links are highly sensitive to atmospheric effects: rapid channel fluctuations induced by turbulence, and slower attenuation due to scattering and absorption [11]. Second, since FSO communication relies on line-of-sight (LOS) alignment with highly directional laser beams and photodetectors, it is vulnerable to pointing errors. These errors can arise from various physical disturbances such as wind-induced vibrations, transmitter-receiver misalignment due to motion, thermal expansion, or structural sway of buildings. In practice, the severity of such errors depends strongly on the acquisition, tracking, and pointing (ATP) subsystem, which mitigates beam misalignment but cannot fully eliminate residual jitter [11]. Either or both of these impairments can significantly degrade the performance and reliability of FSO systems [12].
Several statistical models for the small-scale intensity fluctu- ations are available in the FSO communication literature with a strong basis in realistic terrestrial propagation conditions. We mention a select few important and relevant models as follows: scintillation in weak turbulence is well modeled using log-normal distribution [13], [14], while the scintillation statistic in the weak-to-strong turbulence regimes is best described by gamma-gamma (GG) distribution [14]. Pointing errors due to jitter are addressed in [15] and extended to nonzero boresight in [16]. More recently, generalized turbulence models such as Málaga ($\mathcal{M}$[17], double-generalized gamma [18], and Fisher–Snedecor $\mathcal{F}$ [19] have been introduced, offering enhanced flexibility across diverse channel conditions.
@@
Transmit optical power in terrestrial FSO communications is limited by stringent eye and skin safety regulations [20]. In contrast, FSO links benefit from virtually unlimited optical bandwidths (tens to hundreds of gigahertz per wavelength) to support ultra-high-speed communications [12], making the available power per hertz a potential limiting factor. While such low-SNR operating points are rarely encountered today, recent demonstrations of extreme-wideband wavelength-division multiplexing (WDM) coherent FSO systems indicate that distributing a fixed transmit power over many carriers may plausibly lead to lower per-channel SNR [7], [21]. In particular, variability in atmospheric attenuation under haze, fog, or pollution can accentuate this effect. Thus, operation in the low-SNR regime may emerge as a relevant consideration for future ultra-wideband WDM terrestrial FSO systems. In addition, irradiance fluctuations (mentioned earlier) present in terrestrial FSO channels degrade the power collected at the receiver [11], but this degradation can be alleviated through power control. For instance, a simple scheme in which the transmitter activates only when the instantaneous channel quality exceeds a predefined threshold can be effective under low-power budgets. Importantly, typical terrestrial FSO channels exhibit extremely high coherence bandwidths (hundreds of gigahertz or more) [22, 23] and relatively long coherence times on the order of milliseconds [24]. These conditions yield a quasi-static, frequency-flat block-fading channel, which is increasingly exploited by recent field trials [6, 7, 8, 21] for adaptive transmission in terrestrial coherent FSO links. This context motivates the subsequent analysis of spectral efficiency in low-SNR regimes, where turbulence-induced fluctuations and limited per-Hz transmit power, especially in ultra-wideband WDM coherent FSO systems, can meaningfully affect achievable throughput.
@@
In contrast to the recent practical advances mentioned above, adaptive terrestrial FSO communication systems have attracted comparatively limited rigorous theoretical analysis in the literature, as summarized next. Prior works on IM-DD and coherent HD based terrestrial FSO communication systems have examined achievable spectral efficiency (SE) under log-normal and gamma-gamma (GG) turbulence models (e.g., [25, 26, 27, 28, 29, 30], and references therein). More recent studies have extended this analysis to generalized turbulence channels, such as those modeled by the double-generalized gamma and Málaga ($\mathcal{M}$) distributions [31], [32], [33], [34]. Despite these contributions, the critical aspect of beam power adaptation at the optical transmitter has been largely overlooked—except in a few cases, such as [25], [26], and [27]. This omission is likely due to the analytical complexity involved in solving the associated optimization problems. In a closely related work [26],the SE of an HD-based adaptive FSO system over GG turbulence is examined. However, this analysis does not account for the critical influence of pointing errors, which are particularly significant in terrestrial coherent optical communication systems. Furthermore, [32] presents a unified SE expression for IM-DD and HD based FSO communication systems operating under the $\mathcal{M}$ distribution, which subsumes the GG distribution as a special case. However, this analysis assumes a fixed transmit beam power, without adaptation. It is also important to note that several FSO related studies [25], [28], [29], [31], [32] lack clarity in the methodology used to simulate varying turbulence conditions. In terrestrial FSO systems, several channel parameters are interdependent—for example, changes in the FSO link length simultaneously affect turbulence severity level, path loss, and pointing error statistics. A rigorous understanding of these interrelations grounded in physical modeling is crucial for accurately evaluating FSO system performance. Neglecting such dependencies may lead to misleading interpretations and flawed insights.
To address these research gaps in terrestrial FSO research, we will utilize adaptive optical signal transmission on existing terrestrial FSO communication systems with an emphasis on spectral efficiency improvement, which basically relies on exploiting the temporal channel fluctuations due to turbulence and pointing error impairments. Specifically, this work aims to provide in-depth insights into the operation of terrestrial FSO communication systems, and provide a baseline SE performance for future adaptive coherent FSO systems with strong practical constraints. The principal contributions of this paper are as follows:
  • We present a novel and exact closed-form expression for the average spectral efficiency (ASE) of beam power adaptation-based coherent FSO communication systems with HD receivers, accounting for gamma-gamma (GG) turbulence and pointing error impairments. Theorem 1, along with its specialization in Corollary 1 for the case without pointing errors, represents the central analytical contribution of this work.
  • The accuracy of the derived closed-form ASE expressions is firmly established through extensive Monte Carlo simulations, demonstrating consistent agreement over wide and practically relevant ranges of signal-to-noise ratios (SNRs), turbulence strength, and pointing error conditions, as illustrated in Figures 1-2.
@@
  • To provide deeper insight into the influence of turbulence and pointing errors on adaptive FSO system performance, the exact ASE is further examined in the asymptotic regimes of both high and low SNR regimes. Theorem 2 characterizes these limiting behaviors in detail, thereby providing clear insights into the performance gains achievable as well as the penalties incurred under practical terrestrial FSO links.
  • Among the several concrete findings, two particularly noteworthy and counterintuitive results are highlighted: i) at high SNRs, the ASE of the optical channel improves with turbulence when strong pointing error condition is considered (cf. Fig. 2); and ii) at low SNRs, the ASE of the terrestrial FSO channel improves with turbulence (cf. Fig. 5). These phenomena are rigorously characterized and analytically justified.
@@
The remainder of this paper is organized as follows. The optical turbulence and pointing error models adopted for terrestrial FSO channel are presented in Subsection II-A. The relevant noise source and the coherent FSO system model are described in Subsections II-B and II-C, respectively. Theorems on the spectral efficiency of the turbulent optical channel are derived in Section III and their extensive numerical analysis is presented in Section IV. Finally, Section V summarizes the main results and provides concluding remarks along with directions for future work.
@@ II Channel and System Models
II-A Gamma–Gamma Turbulence with Pointing Errors
The overall intensity fluctuation, or fading, $I$, of a transmitted laser beam due to atmospheric turbulence and pointing error is modeled as
@@ II-A Gamma–Gamma Turbulence with Pointing Errors
Gamma-Gamma Turbulence: The statistics of the intensity fluctuation $I_{\mathrm{a}}$, which spans the weak-to-strong turbulence regimes, are well modeled by the gamma-gamma distribution:
@@
where $a$ and $b$ are shape parameters [14], $\Gamma(\cdot)$ is the Gamma function [36, Eq. (8.310.1)], and $K_{v}(\cdot)$ is the modified Bessel function of the second kind and order $v$ [36, Eq. (8.494.1)]. For plane-wave optical radiation at the receiver, the GG shaping parameters depend on atmospheric conditions as
$a\,$ $=\,\bigl[{\exp\bigl({{0.49\sigma_{\mathrm{\scriptscriptstyle R}}^{2}}}{\bigl({1+1.11\sigma_{\mathrm{\scriptscriptstyle R}}^{2.4}}\bigr)}^{-7/6}\bigr)-1}\bigr]^{-1},$ (3)
$b\,$ $=\,\bigl[{\exp\bigl({0.51\sigma_{\mathrm{\scriptscriptstyle R}}^{2}\bigl(1+0.69\sigma_{\mathrm{\scriptscriptstyle R}}^{2.4}\bigr)^{-5/6}}\bigr)-1}\bigr]^{-1},$ (4)
where $\sigma_{{\mathrm{\scriptscriptstyle R}}}^{2}=1.23\,C_{\mathrm{n}}^{2}k_{\mathrm{w}}^{7/6}{L^{11/6}}$ is the Rytov variance determining the atmospheric turbulence strength, $k_{\mathrm{w}}=2\pi/{\lambda_{\mathrm{w}}}$ is the optical wave number, $\lambda_{\mathrm{w}}$} is the wavelength, $L$ is the propagation distance, and $C_{\mathrm{n}}^{2}$ is the index of refraction structure parameter [14].
@@
where $\text{erf}(\cdot)$ is the standard error function [36, Eq. (8.250.1)], and $0\leqslant A_{\mathrm{\scriptscriptstyle 0}}\leqslant 1$ [15]. For a coherent beam, the beam radius $w_{\mathrm{\scriptscriptstyle L}}$ is related to the beam waist $w_{\mathrm{\scriptscriptstyle 0}}$ at the optical transmitter’s exit aperture by $w_{\mathrm{\scriptscriptstyle L}}=w_{\mathrm{\scriptscriptstyle 0}}(1+\epsilon(\lambda_{\mathrm{w}}L/\pi w_{\mathrm{\scriptscriptstyle 0}}^{2})^{2})^{1/2}$ where $\epsilon=(1+2w_{\mathrm{\scriptscriptstyle 0}}^{2}/\rho_{\scriptscriptstyle 0}^{2}(L))$ is the global coherence parameter, $\rho_{\scriptscriptstyle 0}(L)=(1.46C_{\mathrm{n}}^{2}k_{\mathrm{w}}^{2}L)^{-3/5}$ is the coherence length in turbulence [38], and the transmit beam waist $w_{\scriptscriptstyle 0}$ is related to the transmitter aperture diameter $D$ as $w_{\mathrm{\scriptscriptstyle 0}}={D}/{\sqrt{2}\pi}$ [39]. Farid and Hranilovic [15] derived the PDF of the irradiance component $I_{\mathrm{p}}$ under the assumptions of i.i.d. zero-mean Gaussian distributions for both horizontal and vertical receiver sways (i.e., zero boresight and identical jitters), expressed as
@@
A recent refinement of the pointing error model is introduced in [40], where multiple alternative parameterizations of $\xi$ and $A_{\scriptscriptstyle 0}$ are proposed to enhance modeling accuracy and computational efficiency (see [40, Fig. 5]). In this work, we hence borrow the modified intensity uniform model from [40], which retains the same mathematical form for the probability distribution as in (6), but with redefined parameters given by
@@
Precisely, this choice demonstrates more than an order-of-magnitude reduction in normalized mean square error (NMSE) compared to the Farid–Hranilovic formulation (see [40, Table 1]), thereby offering a substantially more reliable characterization of pointing jitter effects in terrestrial FSO links.
Composite Fading Distribution: Given the distributions in (2) and (6) of the independent constituents $I_{\mathrm{a}}$ and $I_{\mathrm{p}}$, the PDF of the instantaneous composite irradiance $I=I_{\mathrm{a}}I_{\mathrm{p}}$ is developed to [25, Eq. (8)]

### v10

# Maximum Spectral Efficiency of Adaptive Coherent Terrestrial FSO Links

@@ Abstract
Terrestrial free-space optical (FSO) communication systems, while designed to operate on large unlicensed optical bandwidths, are fundamentally power-constrained due to stringent eye-safety regulations. Moreover, channel fluctuations inherent to terrestrial FSO links further reduce the received optical power. Consequently, the achievable signal-to-noise ratio (SNR) per hertz could become limited—especially for future terrestrial FSO systems based on coherent communications. This necessitates the development of efficient and adaptive communication strategies at both the optical transmitter and receiver. However, a comprehensive assessment of adaptive coherent terrestrial FSO systems remains largely unexplored in the existing literature. This paper investigates terrestrial FSO communication links employing adaptive coherent transmission and synchronous heterodyne detection-based reception operating in the shot-noise-limited regime. Specifically, we propose a novel exact closed-form expression for the average spectral efficiency limit of coherent FSO communication systems with optimal adaptive signaling over gamma-gamma turbulence channels with pointing errors. More importantly, we provide a detailed assessment of the impact of turbulence and pointing error impairments on the coherent FSO system performance, revealing several novel and counterintuitive insights. Extensive numerical results help elucidate the intricacies of analyzing these terrestrial FSO systems and clarify a few misconceptions alluded to in recent related literature.
@@ I Introduction
Terrestrial free-space optical (FSO) communication has established itself as a complementary—and in some deployment scenarios, a viable alternative—solution to conventional RF (radio-frequency) and fiber-optic communication technologies for delivering ultra-high-capacity fronthaul and backhaul connectivity. This growth is driven by several key advantages of FSO systems, including vast unlicensed bandwidth, low-cost transceiver implementation, rapid deployment capability, and inherent physical-layer security [1]. Consequently, the scope of FSO-based communication technologies has steadily broadened and now encompasses a diverse range of application scenarios, including wireless ad hoc networks for tactical and emergency response situations [1], terrestrial last-mile broadband networks [2], deep-space communications [3], inter-satellite communication links [4], and mission-critical civil and military operations [5].
A fundamental design consideration in FSO systems concerns how the optical field is exploited for information transfer. Accordingly, two principal transmission and detection paradigms can be identified, with coherent approaches—most commonly realized via heterodyne reception—gradually gaining prominence alongside conventional noncoherent schemes:
@@
(i) Intensity modulation with direct detection (IM-DD): In this scheme, information is encoded solely in the intensity of the transmitted optical field, thereby utilizing only one degree of freedom. At the receiver, the optical field is detected by direct (square-law) photodetection without phase recovery. IM-DD systems are widely adopted in commercial FSO deployments due to their simple low-cost transceiver architectures; however, the limited sensitivity of IM-DD receivers restricts their applicability to short-range terrestrial FSO links [6], [7].
@@
(ii) Coherent (field) modulation with synchronous heterodyne detection (HD): Coherent FSO communication systems exploit both the amplitude and phase of the transmitted optical field, thereby leveraging two degrees of freedom to convey information. In an HD receiver, a local optical oscillator—typically frequency-offset from the received carrier—is mixed with the incoming optical field prior to photodetection. This process enhances receiver sensitivity and preserves the full complex envelope of the received optical field. Moreover, coherent HD receivers enable highly selective narrowband RF filtering, in contrast to the inherently wideband optical filtering in IM-DD FSO systems. Collectively, these capabilities position coherent communication techniques as strong candidates for next-generation high-capacity terrestrial FSO links [7].
Recent advances in coherent optical transceiver techniques tailored to terrestrial FSO links have renewed research interest, with field trials demonstrating per-channel data rates beyond 100 Gbps [8], 400 Gbps [9], and 800 Gbps [10] (see also references therein). While these demonstrations clearly validate the feasibility of coherent FSO communications in terrestrial environments, they also highlight the nontrivial signal processing burden and practical design constraints imposed by the turbulent optical medium, as noted by Guiomar et al. in [11]. Chief among these challenges are channel impairments intrinsic to terrestrial FSO links: turbulence-induced fading and pointing (misalignment) errors. Turbulence-induced refractive-index inhomogeneities in the atmosphere cause random intensity (irradiance) fluctuations, which are typically characterized by statistical channel models such as the lognormal, gamma-gamma (GG), and Málaga distributions, depending on the prevailing turbulence regime [12]. Misalignment induced by structural motion, mechanical disturbances, and strong winds impose additional power penalties, thereby degrading the performance of coherent terrestrial FSO systems. In practice, the severity of pointing errors depends strongly on the acquisition, tracking, and pointing (ATP) subsystem, which mitigates beam misalignment but cannot fully eliminate residual jitter [12].
@@
Beyond turbulence-induced and pointing-induced irradiance fluctuations, optical power emission in terrestrial FSO links is constrained by stringent eye and skin safety regulations [13]. When coupled with the vast optical bandwidths available for ultra-high-speed transmission [14], stringent power limits can render the available power per hertz a dominant performance bottleneck. While such low-SNR operation is rarely encountered today, recent demonstrations of extreme-wideband wavelength-division multiplexed (WDM) coherent FSO systems indicate that allocating the power-constrained optical budget across many optical carriers can plausibly lead to reduced per-channel SNR [15], [16], particularly under variable atmospheric attenuation (e.g., haze, fog, or pollution). Accordingly, the low-SNR regime is likely to gain practical relevance in future ultra-wideband WDM terrestrial FSO links. Moreover, irradiance fluctuations further degrade the received optical power [12]; however, this effect can be alleviated through adaptive transmission strategies. For example, threshold-based transmission—where the transmitter is activated only when the instantaneous channel quality exceeds a predefined level—can be especially effective under stringent power constraints. Importantly, terrestrial FSO links exhibit extremely large coherence bandwidths (hundreds of gigahertz or more) [17, 18], and millisecond-scale coherence times [19], thereby enabling a quasi-static, frequency-flat block-fading channel. Such a block-fading channel structure facilitates adaptive coherent transmission, as demonstrated in recent field experiments [9, 10, 15, 16].
@@
Despite the recent practical advances of adaptive coherent terrestrial FSO communication systems highlighted above, such systems have attracted comparatively limited rigorous theoretical analyses in the literature, as summarized next. Prior works on IM-DD and coherent HD based terrestrial FSO communication systems have examined limits on the achievable average spectral efficiency (ASE) under lognormal and GG turbulence models (e.g., [20, 21, 22, 23, 24, 25], and references therein). More recent studies have extended this analysis to generalized turbulence channels, such as those modeled by the double-generalized gamma and Málaga distributions [26], [27], [28], [29]. Despite these contributions, the crucial aspect of transmitter adaptation has largely been overlooked, with only a few notable exceptions [20], [21], [22]. Among these, [21] investigates a closely related problem, analyzing the ASE of an HD-based adaptive coherent FSO system operating over GG turbulence. However, this analysis neglects the impact of pointing errors, which constitutes a major impairment in practical terrestrial FSO links employing coherent detection. Further, [27] derives a unified ASE expression for IM-DD and HD based coherent FSO communication systems operating under the Málaga distribution—subsuming the GG model as a special case—but assumes no transmitter adaptation. Beyond these analytical limitations, several prior studies [20, 23, 24, 26, 27] provide limited clarity regarding the methodology used to model and simulate varying turbulence conditions. In practical terrestrial FSO deployments, key channel parameters are inherently interdependent; for example, variations in link distance simultaneously influence turbulence strength, path loss, and pointing error statistics. A rigorous understanding of these interrelations, grounded in physically consistent channel modeling, is therefore essential for reliable performance evaluation. Neglecting such dependencies may lead to misleading interpretations and flawed insights. Collectively, these observations highlight the absence of a comprehensive baseline spectral efficiency characterization for adaptive coherent terrestrial FSO systems operating under realistic turbulence conditions, pointing errors, and stringent transmit optical power constraints.
Motivated by the above observations, this paper investigates adaptive transmission for terrestrial coherent FSO links with a focus on spectral efficiency optimization under realistic channel impairments. By explicitly exploiting the temporal fluctuations induced by atmospheric turbulence and pointing errors, the proposed analysis aims to characterize the fundamental performance limits of adaptive coherent FSO systems operating under stringent transmit power constraints. The results provide a baseline spectral efficiency characterization that can serve as a reference for the design and evaluation of future adaptive coherent FSO systems subject to practical deployment considerations. The principal contributions of this paper are summarized as follows:
  • We present a novel and exact closed-form expression for the ASE limit of adaptive coherent transmission with synchronous heterodyne (HD) reception over terrestrial FSO links impaired by GG turbulence and pointing jitters. Theorem 1, together with its specialization in Corollary 1 for the case without pointing errors, constitutes the primary analytical contribution of this work.
@@
  • To provide deeper insights into the impact of turbulence and pointing errors on adaptive FSO system performance, the exact ASE limit is further analyzed in the asymptotic high-SNR and low-SNR regimes. Theorem 2 rigorously characterizes these limiting behaviors, thereby elucidating both the achievable performance gains and the inherent penalties under practical terrestrial FSO conditions.
  • The validity of the derived ASE limit expressions is rigorously verified through extensive Monte Carlo simulations, which demonstrate excellent agreement across wide and practically relevant ranges of SNR, turbulence strength, and pointing error, as illustrated in Figures 12.
  • Among the key findings, two particularly noteworthy and counterintuitive insights are identified: i) at high SNRs, the relative ASE degradation diminishes as turbulence strengthens under severe pointing-error conditions (cf. Fig. 2); and ii) at low SNRs, the ASE increases as turbulence strengthens (cf. Fig. 5). These effects are rigorously characterized and analytically substantiated.
@@
Paper Organization: The optical turbulence and pointing error models adopted for terrestrial FSO channels are presented in Subsection II-A. The relevant noise source and the adaptive coherent FSO system model are described in Subsections II-B and II-C, respectively. Theorems on the ASE limit of the terrestrial coherent FSO channel are derived in Section III, and their extensive numerical analysis is presented in Section IV. Finally, Section V summarizes the main results and presents concluding remarks and future research directions.
@@ II Channel and System Models
II-A Optical Turbulence with Pointing Errors
The overall intensity fluctuation, or fading, $I$, of a transmitted laser beam due to atmospheric turbulence and pointing errors is modeled as
@@ II-A Optical Turbulence with Pointing Errors
Gamma-Gamma (GG) Turbulence: The statistics of the intensity fluctuation $I_{\mathrm{a}}$ across weak-to-strong turbulence regimes are well modeled by the GG distribution, given by
@@
where $a$ and $b$ are shape parameters [31], $\Gamma(\cdot)$ is the Gamma function [32, Eq. (8.310.1)], and $K_{v}(\cdot)$ is the modified Bessel function of the second kind and order $v$ [32, Eq. (8.494.1)]. For plane-wave optical radiation at the receiver, the GG shaping parameters depend on atmospheric conditions as
$a\,$ $=\,\biggl[\,{\exp\,\biggl({{0.49\sigma_{\mathrm{\scriptscriptstyle R}}^{2}}\,}{\Bigl({1+1.11\sigma_{\mathrm{\scriptscriptstyle R}}^{12/5}}\Bigr)}^{-7/6}\biggr)-1}\,\biggr]^{-1},$ (3)
$b\,$ $=\,\biggl[\,{\exp\,\biggl({0.51\sigma_{\mathrm{\scriptscriptstyle R}}^{2}\,\Bigl(1+0.69\sigma_{\mathrm{\scriptscriptstyle R}}^{12/5}\Bigr)^{-5/6}}\biggr)-1}\,\biggr]^{-1},$ (4)
where $\sigma_{{\mathrm{\scriptscriptstyle R}}}^{2}=1.23\,C_{\mathrm{n}}^{2}k_{\mathrm{w}}^{7/6}{L^{11/6}}$ is the Rytov variance determining the atmospheric turbulence strength, $k_{\mathrm{w}}=2\pi/{\lambda_{\mathrm{w}}}$ is the optical wave number, $\lambda_{\mathrm{w}}$} is the wavelength, $L$ is the propagation distance, and $C_{\mathrm{n}}^{2}$ is the index-of-refraction structure parameter [31].
@@
where $\text{erf}(\cdot)$ is the standard error function [32, Eq. (8.250.1)], and $0\leqslant A_{\mathrm{\scriptscriptstyle 0}}\leqslant 1$ [34]. For a coherent beam, the beam radius $w_{\mathrm{\scriptscriptstyle L}}$ is related to the beam waist $w_{\mathrm{\scriptscriptstyle 0}}$ at the optical transmitter’s exit aperture by $w_{\mathrm{\scriptscriptstyle L}}=w_{\mathrm{\scriptscriptstyle 0}}(1+\epsilon(\lambda_{\mathrm{w}}L/\pi w_{\mathrm{\scriptscriptstyle 0}}^{2})^{2})^{1/2}$ where $\epsilon=(1+2w_{\mathrm{\scriptscriptstyle 0}}^{2}/\rho_{\scriptscriptstyle 0}^{2}(L))$ is the global coherence parameter, $\rho_{\scriptscriptstyle 0}(L)=(1.46C_{\mathrm{n}}^{2}k_{\mathrm{w}}^{2}L)^{-3/5}$ is the coherence length in turbulence [35], and the transmit beam waist $w_{\scriptscriptstyle 0}$ is related to the transmitter aperture diameter $D$ as $w_{\mathrm{\scriptscriptstyle 0}}={D}/{\sqrt{2}\pi}$ [36]. Farid and Hranilovic [34] derived the PDF of the irradiance component $I_{\mathrm{p}}$ under the assumptions of i.i.d. zero-mean Gaussian distributions for both horizontal and vertical receiver sways (i.e., zero boresight and identical jitters), expressed as
@@
A recent refinement of the pointing error model is introduced in [37], where multiple alternative parameterizations of $\xi$ and $A_{\scriptscriptstyle 0}$ are proposed to enhance modeling accuracy and computational efficiency (see [37, Fig. 5]). In this work, we hence borrow the modified intensity uniform model from [37], which retains the same mathematical form for the probability distribution as in (6), but with redefined parameters given by
@@
Precisely, this choice demonstrates more than an order-of-magnitude reduction in normalized mean square error (NMSE) compared to the Farid–Hranilovic formulation (see [37, Table 1]), thereby offering a substantially more reliable characterization of pointing jitter effects in terrestrial FSO links.
Composite Fading Distribution: Given the distributions in (2) and (6) of the independent constituents $I_{\mathrm{a}}$ and $I_{\mathrm{p}}$, the PDF of the instantaneous composite irradiance $I=I_{\mathrm{a}}I_{\mathrm{p}}$ is developed to [20, Eq. (8)]